Tohban Report 2011-01-05

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 29 Dec 2010
End Date: 05 Jan 2011
Tohban: Pascal Saint-Hilaire
Tohban email:
Next Tohban: Jim McTiernan
List all reports


Solar Activity

Small ARs on the Sun have produced mild activity this week: 21 B-flares, 2 C-flares.

Memory Management

SSR reaches 0% during last or penultimate of the daily pass set. Max is 25%. Given the SSR's good behaviour, it has been decided that we will turn on the *rear* segments during RHESSI night.

Data Gaps

Glitch yesterday at Wallops - data will be replayed. Otherwise, only one 300 s long VC3 gap on Monday.

Spacecraft Status

Spin rate is 15 rpm and the cold plate 2 temperature is 100 K, normal due to precession. But there is an overall upward trend since last two years. 100K is the (probably rule-of-thumbish) threshold for extra-vulnerability to radiation damage. Brian Dennis will send an email to D. Smith asking whether the cryocooler power (currently 71 W) should be increased a little bit. Ops think that should be ok, but to be verified.

Personal tools